SUMMARY OF LEG 182 ENGINEERING AND DRILLING OPERATIONS
Transit to Site 1132
The 44-nmi sea voyage to Site 1132 required 4.25 hr at 10.4 kt. A beacon was deployed at 2221 hr on 20 November, initiating Site 1132.
The ship was stabilized on position, and Hole 1132A was spudded at 0100 hr on 21 November. The bit was positioned at 229.5 m, and Core 1H recovered 9.22 m. However, the excessive recovery was not appropriate for the establishment of a good mudline and the hole was terminated.
The ship was not moved and the bit was positioned at 227.5 m, and Hole 1132B was spudded at 0130 hr on 21 November. Core 1H recovered 6.78 m, indicating a water depth of 218.5 mbsl, which suggests the previous assessment of an inappropriate mudline in Hole 1132A was correct. APC coring advanced to 168.3 mbsf, ceasing after four of the five last APC cores were recovered with split liners. Cores 3H18H were oriented and Adara tool heat-flow measurements were taken on Cores 4H, 8H, and 12H. The nonmagnetic shoe and flapper core catcher with a steel 10-finger core catcher were run on Cores 3H, 5H, 7H, 9H, 11H, and 13H. A DVTP temperature probe was deployed after Core 18H at 168.3 mbsf. XCB cores deepened the hole from 168.3 to 284.6 mbsf. Drilling times increased after Core 24X (219.5 mbsf) and recovery was poor below Core 28X (257 mbsf), thus Hole 1132B was terminated four cores later. The fishing vessel Bolzano was hired to deliver to the Resolution six self-contained breathing apparatus, two air-tank filler fittings, and 10 air-filling station filters. The Bolzano arrived at Site 1132 at 0200 hr on 21 November, discharged its cargo in four crane lifts, and departed at 0315 hr. The hole was plugged with mud, and the drill string was pulled to the rig floor at 0425 hr on 22 November, ending Hole 1132B.
The ship was moved 20 m northwest, and the standard RCB/BHA was run to the seafloor, spudding Hole 1132C at 0735 hr on 22 November. The hole was drilled to 161.6 mbsf with a center bit in 3.5 hr. The center bit was retrieved and RCB Cores 1R and 2R were cut from 161.6 to 180.4 mbsf in a successful effort to sample a low recovery interval from Hole 1132B. The hole was then drilled from 180.4 to 255.8 mbsf in 3.25 hr. Coring resumed with the RCB from 255.8 to 603.2 mbsf with low recovery in interbedded cherts and packstones. With no recovery in the last four cores and vigorous backflow occurring on pipe connections, a precautionary wiper trip was made to verify hole conditions. The pipe was pulled up to 535 mbsf with heavy backflow. The bit deplugger was dropped in case the annulus was plugged with debris, causing the backflow. The bottom was tagged at 574 mbsf, indicating 29 m of fill in the bottom of the hole. The hole was cleaned to continue coring, ultimately clearing out 9 m of the fill. However, the pipe became stuck with 40 kips overpull, stalling the rotary. The pipe was worked free, and another mud sweep was circulated. A full wiper trip was made to 105 mbsf with heavy backflow and fill was tagged at 560 mbsf (43 m off the bottom). With these unstable conditions, it appeared unlikely that the hole could be safely deepened at these shallow-water depths. Thus, coring was terminated and another mud sweep was circulated, the bit was dropped with the MBR, and the sleeve was closed. The hole was loaded with 150 bbl of mud in an effort to stabilize the hole.
The end of pipe was pulled to 105 mbsf for logging. The Triple combo log was run from 470 to 397 mbsf and rerun from 560 mbsf to the mud line. The FMS-Sonic tool could not be worked past 163 mbsf, so it was run from 163 to 80 mbsf. The open-ended pipe was advanced from 105 to 89 mbsf to guide the logging tool through a presumed ledge; however, the FMS Sonic tool could not pass 203 mbsf. The pipe was advanced again to 242 mbsf to push past any obstructions, where it stopped with 25 kips weight on the pipe. The drill string started hydraulically lifting when an attempt was made to circulate, which suggested that it was buried in fill. The end of pipe was pulled up to 105 mbsf and the FMS-Sonic tool was rerun in attempt to log the upper portion of the hole. However, the tool could not pass 180 mbsf, and the log was run from that depth. The hole was plugged with mud, the drill string retrieved, and both beacons recovered. The rig was secured for transit and the vessel was under way at 1815 hr on 24 November.
To 182 Summary of Engineering and Drilling Operations for Site 1133
To 182 Table of Contents